# suppSupplementary References

### v1

@@ 4 Discussion and Future Work
Limitations. Our evaluation demonstrates that Rawsamble achieves high throughput, making it a viable candidate for real-time analysis while constructing de novo assemblies. However, there are still two main challenges to fully utilize real-time de novo assembly construction during sequencing. First, the hash-based index should be constructed and updated dynamically in real-time while sequencing is in progress. Although Rawsamble provides the mechanisms for storing multiple hash tables that can be constructed for each chunk of raw signals generated in real-time, it is not computationally feasible to dynamically update the index for all sequenced signals as the memory and computational burden increases with each hash table generated. Such an approach requires a decision-making mechanism to dynamically stop updating the index after sequencing a certain amount of signals. The stopping mechanism should be accurate enough to ensure that the current state of the index provides sufficient information to find the overlaps between the already sequenced signals and the new signals generated after the index construction. Second, the assembly graph should be constructed and updated dynamically while the new overlap information is generated in real time. This is challenging as the intermediate steps for generating the unitigs (e.g., the transitive reduction step [87]) in assembly graphs may not work optimally without the full overlap information, as it is likely to remove graph connections that can be useful with new overlap information. It might be feasible to use graph structures that are more suitable for streaming data generation, such as de Bruijn graphs, for assembly construction purposes [88, 89, 90].
@@
We evaluate Rawsamble using raw signals generated from ONT’s R9.4 flowcells to show that overlapping information between raw signals can be used to construct de novo assemblies. We leave optimizations for newer flow cell versions (e.g., R10.4.1) for future work.
@@ 5 Conclusion
We introduce Rawsamble, the first mechanism that can find overlaps between two sets of raw nanopore signals without translating them to bases. We find that Rawsamble can 1) find overlaps while meeting the real-time requirements with an average throughput of $2{,}432{,}561$ signals/sec per CPU thread, 2) reduce the overall time needed for finding overlaps (on average by 16.36×16.36$\times$16.36 × and up to by 41.59×41.59$\times$41.59 ×) and peak memory usage (on average by 11.73×11.73$\times$11.73 × and up to by 41.99×41.$9$) compared to the time and memory needed to run state-of-the-art read mapper (minimap2) combined with the Dorado basecaller running on a CPU with its fastest model, 3) share a large portion of overlapping pairs with minimap2 ($36.57\%$ on average), and 4) construct long assemblies from these useful overlaps. We find that we can construct assemblies of half the length of the entire E. coli genome (i.e., of length 2.7 million bases) directly from raw signal overlaps without basecalling. Finding overlapping pairs from raw signals is critical for enabling new directions that have not been explored before for raw signal analysis, such as de novo assembly construction from overlaps that we explore in this work. We discuss many other new directions that can be enabled by finding overlaps and constructing de novo assemblies.
@@
Acknowledgments
SAFARI Research Group acknowledges the generous gift funding provided by our industrial partners (especially by Google, Huawei, Intel, Microsoft, VMware), which has been instrumental in enabling the 15+ year-long research SAFARI Research Group has been conducting on accelerating genome analysis. This work is also partially supported by the Semiconductor Research Corporation (SRC), the European Union’s Horizon programme for research and innovation [101047160 - BioPIM] and the Swiss National Science Foundation (SNSF) [200021_213084]. M.M. has been supported by the Max Planck ETH Center for Learning Systems and by SNF Project Grant #200550 to A.K. S.G. is funded through Swiss National Science Foundation Project grant no. 200550 to A.K.
@@ Acknowledgments
Rawsamble: Overlapping and Assembling Raw Nanopore Signals
@@ A RawHash2 Overview
First, to generate sequences of signals that can be compared to each other, RawHash2 generates signals of k-mers, called events, from both a reference genome and raw signals. To generate events from reference genomes, it uses a lookup table, called k-mer model, that provides the expected signal value (i.e., event value) as a floating value for each possible k-mer where k is usually 6 or 9, depending on the flow cell version. To identify events (i.e., k-mers) in raw signals, RawHash2 performs a segmentation technique to detect the abrupt changes in signals, which enables identifying the regions in signals generated when sequencing a particular k-mer. RawHash2 uses the average value of signals within the same region as an event value after identifying outliers within the region. Due to the variations and noise in nanopore sequencing, event values generated from the same k-mer can slightly differ from each other, making it challenging to directly match the event values to each other to identify matching k-mers between a reference genome and raw signals.
Second, to mitigate this noise issue, RawHash2 quantizes the event values such that slightly different event values can be quantized into the same value (i.e., bucketing) to enable direct matching of quantized event values between a reference genome and raw signals. To enable an accurate quantization, RawHash2 identifies the range of values that are assigned to the same quantized value dynamically according to the nanopore model.
@@
B Visualizing Assemblies
[…]
Figure S2: Visualization of the assembly graphs generated using Rawsamble and minimap2 overlaps from the D2 (E. coli) dataset.
Figure S3: Visualization of the assembly graphs generated using Rawsamble and minimap2 overlaps from the D3 (Yeast) dataset.
Figure S4: Visualization of the assembly graphs generated using Rawsamble and minimap2 overlaps from the D4 (Green Algae) dataset.
Figure S5: Visualization of the assembly graphs generated using Rawsamble and minimap2 overlaps from the D5 (Human) dataset.
C Generating a Gold Standard Assembly
@@
For the high-coverage D2 E.coli dataset, Hifiasm outputs a fragmented assembly, while Verkko requires extensive parameter tuning (specifically in terms of coverage, ---unitig-abundance, and ---base-k) to achieve desirable contiguity and completeness. LJA produces an almost perfect assembly when a minimum length filter of 30kbp is applied to the corrected reads, as suggested by the HERRO paper. Flye performs comparably to LJA without the need for error correction.
@@
D2 E. coli 445$\times$ 240$\times$
D3 Yeast 32$\times$ 12$\times$
D4 Green algae 5.6$\times$ 3.7$\times$
D5 Human 0.6$\times$ 0.002$\times$
D Configuration
D.1 Parameters
@@
Tool D1 SARS-CoV-2 D2 E. coli D3 Yeast D4 Green Algae D5 Human
Rawsamble -x ava-viral -t 32 -x ava -t 32 -x ava -t 32 -x ava -t 32 -x ava –chain-gap-scale 0.6 -t 32
@@
Dorado CPU (Fast) basecaller -x cpu dna_r9.4.1_e8_fast@v3.4
Dorado CPU (HAC) basecaller -x cpu dna_r9.4.1_e8_hac@v3.3
Dorado GPU (HAC) basecaller dna_r9.4.1_e8_hac@v3.3
Dorado GPU (SUP) basecaller dna_r9.4.1_e8_sup@v3.3
Miniasm
@@ D.2 Versions
Dorado 0.7.3 https://github.com/nanoporetech/dorado/releases/tag/v0.7.3

### v2

@@ 4 Discussion and Future Work
Limitations. Our evaluation demonstrates that Rawsamble achieves high throughput when finding overlaps between raw signal pairs, making it a viable candidate for real-time analysis. However, there are still two main challenges to fully utilize real-time de novo assembly construction during sequencing. First, the hash-based index should be constructed and updated dynamically in real-time while sequencing is in progress. Although Rawsamble provides the mechanisms for storing multiple hash tables that can be constructed for each chunk of raw signals generated in real-time, it is not computationally feasible to dynamically update the index for all sequenced signals as the memory and computational burden increase with each hash table generated. Such an approach requires a decision-making mechanism to dynamically stop updating the index after sequencing a certain amount of signals. The stopping mechanism should be accurate enough to ensure that the current state of the index provides sufficient information to find the overlaps between the already sequenced signals and the new signals generated after the index construction. Second, the assembly graph should be constructed and updated dynamically while the new overlap information is generated in real time. This is challenging as the intermediate steps for generating the unitigs (e.g., the transitive reduction step [83]) in assembly graphs may not work optimally without the full overlap information, as it is likely to remove graph connections that can be useful with new overlap information. It might be feasible to use graph structures that are more suitable for streaming data generation, such as de Bruijn graphs, for assembly construction purposes [84, 85, 86].
@@
Challenges for constructing and evaluating de novo assemblies. Although the main focus of our work is enabling all-vs-all overlapping between raw signals, we identify de novo assembly construction as a natural use case that can utilize the all-vs-all overlap information from raw signals that Rawsamble finds. To show that the overlaps from Rawsamble can lead to unitigs that are much longer than average read lengths, we construct and evaluate the de novo assembly graphs from the Rawsamble overlaps, as discussed in Section 3.4. However, we identify two key challenges that limit the scope of our evaluation of these de novo assemblies, which we leave as future work.
First, we cannot generate the sequence of assembled signals from the assembly graphs (i.e., spelling the sequence from the assembly graph). This is because we use the miniasm assembler to construct these assembly graphs from the raw signal overlaps. Although miniasm can use the overlap position information to construct the assembly graph, it cannot output the sequence of assemblies, as it is mainly designed for basecalled sequences. This is challenging because each raw signal contains metadata used in basecalling mechanisms (e.g., nanopore-specific information such as the baseline current levels). Carrying the metadata information effectively from many overlapping signals is essential to enable further analysis of these assembled signals.
Second, to evaluate the identity of the constructed assemblies, these assemblies need to be aligned to a ground truth assembly. This can be done either by either 1) basecalling the assembled signal or 2) aligning the assembled signal using dynamic time warping (DTW) [57] to the ground truth assembly. These approaches require constructing the assembled signal (i.e., the first challenge mentioned above) and designing a new class of basecallers to directly basecall assembled signals, which we discuss as future work below. Although the missing identity evaluations of these assemblies can be a major limitation of our work, we believe the assemblies that are constructed from the Rawsamble overlaps are likely to be close to the assembly accuracy that can be achieved by using the minimap2 overlaps based on our estimated accuracy evaluation.
@@ 5 Conclusion
We introduce Rawsamble, the first mechanism that can find overlaps between two sets of raw nanopore signals without translating them to bases. We find that Rawsamble can 1) find overlaps while meeting the real-time requirements with an average throughput of $1{,}496{,}148$ signals/sec per CPU thread, 2) reduce the overall time needed for finding overlaps (on average by 4.48×4.48$\times$4.48 × and up to by 23.10×23.10$\times$23.10 ×) and peak memory usage (on average by 4.33×4.$3$ and up to by 22.00×22.$0$) compared to the time and memory needed to run state-of-the-art read mapper (minimap2) combined with the Dorado basecaller running on a CPU with its fastest model, 3) share a large portion of overlapping pairs with minimap2 ($30.94\%$ on average), and 4) construct long assemblies from these useful overlaps. We find that we can construct assemblies of half the length of the entire E. coli genome (i.e., of length 2.3 million bases) directly from raw signal overlaps without basecalling. Finding overlapping pairs from raw signals is critical for enabling new directions that have not been explored before for raw signal analysis, such as de novo assembly construction from overlaps that we explore in this work. We discuss many other new directions that can be enabled by finding overlaps and constructing de novo assemblies.
@@
Rawsamble: Overlapping Raw Nanopore Signals
@@ A RawHash2 Overview
[…]
Second, to further mitigate this noise issue in events, RawHash2 quantizes the event values such that slightly different event values can be quantized into the same value (i.e., bucketing) to enable direct matching of quantized event values between a reference genome and raw signals. To enable an accurate quantization, RawHash2 identifies the range of values that are assigned to the same quantized value dynamically according to the nanopore model.
@@
B Generating a Gold Standard Assembly
@@
For the high-coverage D1 E.coli dataset, Hifiasm outputs a fragmented assembly, while Verkko requires extensive parameter tuning (specifically in terms of coverage, ---unitig-abundance, and ---base-k) to achieve desirable contiguity and completeness. LJA produces an almost perfect assembly when a minimum length filter of 30kbp is applied to the corrected reads, as suggested by the HERRO paper. Flye performs comparably to LJA without the need for error correction. It should be noted that the golden standard assembly for D4 E.coli dataset was limited in terms of assembly coverage, using ---asm-coverage 50, as Flye raised an error regarding an otherwise too high coverage level.
@@
D1 E. coli 445$\times$ 240$\times$
D2 Yeast 32$\times$ 12$\times$
D3 Green algae 5.6$\times$ 3.7$\times$
D4 E. coli (R10.4.1) 1796.9$\times$ 338$\times$
C Configuration
C.1 Parameters
@@
Tool D1 E. coli D2 Yeast D3 Green Algae D4 E. coli (R10.4.1)
Rawsamble -x ava -t 32 -x ava -t 32 -x ava -t 32 -x ava –r10 -t 32
@@
Dorado CPU (Fast) basecaller -x cpu fast
Dorado CPU (HAC) basecaller -x cpu hac
Dorado GPU (HAC) basecaller hac
@@ C.2 Versions
Dorado 0.9.6 (for R9.4) https://github.com/nanoporetech/dorado/releases/tag/v0.9.6
Dorado 1.0.2 (for R10.4.1) https://github.com/nanoporetech/dorado/releases/tag/v1.0.2